### v1

@@ 2.2 The tidal decay of hot Jupiters
Because hot Jupiters are close to their host stars, the tidal interactions can reduce the orbital eccentricity ($e$) of close-in planets [7, 12]. Tides can also transfer orbital energy and angular momentum from the hot Jupiters to their host stars, resulting in the spin-up of the stars [72, 73, 74, 75, 76]. As a consequence, the semi-major axis ($a$) of hot Jupiters would shrink with age. Part of them may have already passed into the Roche limit and been tidally disrupted by the host star [12, 77].
In order to quantitatively determine the tidal reduction of the semi-major axis $a$, we adopt the classical equations of equilibrium tidal evolution: [78, 13]
@@
where $G$ is the gravitational constant, $R$ and $M$ are the radius and mass, and $Q^{{}^{\prime}}$ is the modified tidal dissipation factor [78]. The subscripts * and $\rm p$ refer to the star and planet, respectively. As can been, due to the tides, $a$ decays with the increase of age. The tidal decay rate is dependent on various parameters (e.g., the initial $a$ and $Q^{{}^{\prime}}_{*}$).
Plenty of previous studies have estimated the stellar tidal quality factor $Q^{{}^{\prime}}_{*}$, using various samples and methodologies. Specifically, Jackson et al. (2008) [12], Bonomo et al. (2017) [17], Penev et al. (2018) [24] and Labadie-Bartz et al. (2019) [79] provided constraints based on the evolution of planetary semi-major axis ($a$), eccentricity ($e$), and stellar spin-up. Whereas, Meibom et al. (2005) [40] and Milliman et al. (2014) [41] restrict the $Q^{{}^{\prime}}_{*}$ from the tidal interaction between stellar binaries. Collier Cameron & Jardine (2018) [18]estimated the tidal quality factor by reproducing the observed orbital separation distribution, assuming an initially uniform distribution inferred from Doppler surveys. Hamer & schlaufman (2019) [19] obtains a upper limit of $\log Q^{{}^{\prime}}_{*}\lesssim 10^{7}$ by assuming that hot Jupiters were destroyed before the end of the main sequence lifetime of their host stars. Recently, Yee et al. (2020) [14] and Turner et al. (2021) [15] constrained $Q^{{}^{\prime}}{*}$ by analyzing the orbital decay of WASP-12b, a transiting hot Jupiter with a 1.09-day orbit around a late-F star. Additionally, some other studies have detected hints of orbital decay for tens of hot Jupiters (e.g., WASP-4b, 43b) and provided estimations for $Q^{{}^{\prime}}_{*}$ [80, 81]. We have summarized the constraints on the value of $Q^{{}^{\prime}}_{*}$ from the literature in Extended Data Tab. 1.
@@ 2.3 Numerical simulation
    ‘Late’ model. The formation timescale obeys a distribution referring to Fig. 9 of Hamers et al. (2017) [31].
@@
For the subsequent tidal evolution, we use a simplified model to perform numerical simulations by setting the orbital eccentricity $e=0$ since the circularization timescale is order of magnitudes less than the in-sprial timescale for hot Jupiters [12, 13]. That is to say, the conditions for a hot Jupiter’s ‘arrival’ defined in this work refer to it entering within 0.1 AU of its star, with its orbit having been largely circularized. The simulations are carried out by adopting the eighth-order Runge-Kutta method and the precision is set as $10^{-12}$. As shown in Equation S13, the initial conditions consist of the stellar tidal dissipation factor $Q^{{}^{\prime}}_{*}$, the masses and radii of star and planet, and the initial period distribution of hot Jupiters.
@@
    ‘Early’ model: For the planetary mass and radius, we adopt the same distributions as the observed warm/cold Jupiters. For the initial orbital period distribution, we here adopt a distribution inferred from the confirmed planets in Kepler field [82].
@@
    ‘Late’ model: mass and period as the results of Fig. 8 and 3 of Hamers et al. (2017) [31]. The planetary radius are estimated with the planetary mass-radius relation, $\frac{R_{\rm p}}{R_{\rm J}}=1.2\left(\frac{M_{\rm p}}{M_{\rm J}}\right)^{0.27}$ [83].
@@
In each set of simulation, we put 10,000 hot Jupiters with the aforementioned initial conditions. During the simulations, the hot Jupiters were removed from the system when they fell into the Roche limit, $a_{R}=(R_{p}/0.462)(M_{*}/M_{p})^{1/3}$ [84]. $\log Q^{{}^{\prime}}_{*}$ is set at the interval of 0.1 from 4 to 9. Upon completion of the simulations, the returned information consist of the simulation time, semi-major axis, and the lifetimes of each hot Jupiter.
@@ 3.2 Additional evidence from the obliquity distributions
Stellar obliquity refers to the angle between a planet’s orbital axis and its host star’s spin axis, which can provide important clues to the formation and dynamic evolution of hot Jupiters. Specifically, the initial distribution of stellar obliquity is determined by the formation and evolution of a protoplanetary disk [85, 86]. Subsequently, the stellar obliquity can be further excited by the dynamic mechanisms via interactions caused by planetary and/or binary companions, e.g., planet–planet scattering , stellar/planetary Lidov–Kozai (LK) effects secular chaos. The predicted stellar obliquity distributions of the hot Jupiters via different mechanisms are different (see fig. 22 of [87]). For example, the majority of hot Jupiter systems caused by the planet-planet scattering mechanism would be aligned but a small fraction of systems would have a diverse range of stellar obliquities [67]. In contrast, the secular chaos mechanism will result in misaligned systems with obliquities $\lesssim 90^{\circ}$ [34]. Lastly, the tidal interaction could realign the host star’s spin axis with the planet’s orbital axis and thus reduce the stellar obliquity [87]. Therefore, investigating the distribution of the stellar obliquities of hot Jupiter systems can help to constrain their formation channel and tidal evolution [36, 88].
From observations, typically only the sky-projected stellar obliquity $\lambda$ (the projections onto the plane of the sky) can be mainly measured through the Rossiter–McLaughlin effect [89, 90]. To date, the stellar obliquities have been measured for hundreds of hot Jupiters. Previous studies have found observational evidence that many hot Jupiters in misaligned systems are late-arrived within $\lesssim 0.1$ AU long after their parent protoplanetary disks dissipated [36]. Dong & Foreman-Mackey (2023) [88] reported that most hot Jupiter system are aligned and the other misaligned systems have nearly isotropic stellar obliquities with no strong clustering near 90 degrees.
We initialize the sample of hot Jupiters with $\lambda$ measurements based on the Table A of [87]. We also compare with the TEPCat orbital obliquity catalog (https://www.astro.keele.ac.uk/jkt/tepcat/obli- quity.html) and remove hot Jupiters with the differences in $\lambda$ larger than three times of uncertainties. Then we cross-match with our planetary sample and yield 39 hot Jupiters with $\lambda$ measurements.
As mentioned before, the tidal interaction would reduce the observed $\lambda$. The timescale of tidal-realignment is not well determined yet [91]. In the equilibrium tidal theory, the timescale for realignment is comparable to the in-sprial timescale [87], which is dependent on the stellar mass, radius and planetary mass, radius, orbital semi-major axes. Referring to [87], we adopt the dimensionless tidal dissipation parameter,
@@
where $M_{*}$ and $R_{*}$ are the stellar mass and radius. $M_{\rm p}$ and $a$ are the planetary mass and semi-major axes. Then we select hot Jupiter systems with long-alignment timescales using the following criterion: $\left(\frac{M_{\rm p}}{M_{*}}\right)^{-2}\left(\frac{a}{R_{*}}\right)^{6}>10^{12}$ or $T_{\rm eff}>6250$ K [87]. Furthermore, as shown in Fig. 11 of [87], hot Jupiter systems older than $\sim 4$ Gyr have relatively smaller $\lambda$, thus we also only keep hot Jupiter systems with kinematic ages <4<4 Gyr. We also make Kolmogorov–Smirnov (KS) tests and check the resulted $p$ values to quantify the age evolution.
@@
To further put constraints on the origin of hot Jupiters, we then compare the observed $\lambda$ distribution with the predictions of different formation mechanisms, i.e., planet-planet scattering [67], resonance crossing [92], secular chaos [34], and planetary Kozai-Lidov effect [93]. We also conduct an aligned distribution of stellar obliquity $\phi$ as uniformly distributed in the range of $0-20^{\circ}$. Specifically, we first generate 1,000 hot Jupiter systems for each of the above mechanisms with the distributions of stellar obliquity $\phi$. Then we obtain the sky-projected obliquity $\lambda$ distribution from the $\phi$ distribution with the following formula [88]:
@@
where $\theta$ is the azimuthal angle of the stellar spin axis relative to the orbital axis, and $i_{\rm orb}$ is the orbital inclination. Referring to [88], we take $\theta$ and $\cos{i_{\rm orb}}$ as uniformly distributed between 0 and 1. Finally, we perform two-sample KS tests between the $\lambda$ distributions of the observed ‘early-arrived’/‘late-arrived’ hot Jupiter systems and the prediction of the five mechanisms.
The resulted $p$-values are summarized as Extended Data Tab. 3. As can be seen, the $\lambda$ distribution of the ‘early-arrived’ population matches well with the predictions of the planet-planet scattering [67], the resonance cross [92], and the alignment distribution, but is significantly smaller than the prediction of planetary Kozai-Lidov mechanism [93] with a KS $p$-value <0.003<0.003. For the ‘late-arrived’ hot Jupiters, their $\lambda$ distribution is consistent with the prediction of the secular chaos mechanism [34] with a KS $p$-value of 0.229, but it deviates from the predictions of all other mechanisms with KS $p$-values<0.05<0.05. The most-favored mechanisms (with largest $p$-values, highlighted in Extended Data Tab. 3) for the ‘early-arrived’ and ‘late-arrived’ hot Jupiters are planet-planet scattering and secular chaos, respectively.
@@ 3.3 Consistency check of stellar metallicity
Stellar metallicity also has important impacts on the formation and evolution of hot Jupiters [94]. From theory, gentle disk migration is expected to operate in environments with a range of metallicities. Whereas high eccentricity migration via secular chaos occurs in systems with multiple giant planets, which are more likely to form around metal-rich stars born in disks with more solids.
Stellar metallicity is known to be correlated with other stellar properties [95]. Thus we adopt the NearestNeighbors method in scikit-learn to select the nearest neighbor in the stellar mass and $TD/D$ from the ‘early-arrived’ hot Jupiter hosts for every ‘late-arrived’ hot Jupiter host star. As shown in Extended Data Fig. 7, after eliminating the effects of stellar mass, radius and age, the ‘early-arrived’ hot Jupiter host stars have a relative smaller $\rm[Fe/H]$ comparing to the ‘late-arrived’ hot Jupiter hosts. We generate the 10,000 bootstrapped sample from the $\rm[Fe/H]$ of ‘late-arrived’ and ‘early-arrived’ hot Jupiter hosts (after parameter controlling) and calculate their average metallicities $\overline{\rm[Fe/H]}$. The uncertainties are taken as the $50\pm 34.1$ percentile. The $\overline{\rm[Fe/H]}$ are $0.06^{+0.02}_{-0.03}$ and $0.12^{+0.02}_{-0.02}$ dex for the ‘early-arrived’ and ‘late-arrived’ hot Jupiter hosts, respectively. Of the 10,000 set of bootstrapping, the $\overline{\rm[Fe/H]}$ of ‘late-arrived’ hot Jupiter hosts are larger for 9,607 times, supporting our inference that the ‘late-arrived’ hot Jupiters are formed via secular chaos.
@@ 3.4 Consistency check of configurations of hot Jupiter systems
Previous studies have detected hints of the orbital decay of hot Jupiters through monitoring of their transit-time variations (TTVs) [14, 80]. Recently, by combining the high-precision 2 minute cadence transit data provided by TESS with archival data from previous work, Wang et al. (2024) identified 11 candidates with decreasing orbital periods that also passed the leave-one-out cross validation test from 326 hot Jupiters [16]. Out of the 11 candidates, only the orbital decay of WASP 12 b has been confirmed [14, 15]. For the other 10 candidates, the possibility that mechanisms other than tidal interactions (e.g., apsidal precession [96], the Rømer effect [97], and the Applegate effect[98]) may contribute to the variations in orbital periods cannot be ruled out. Therefore, 1 and 11 can be taken as the lower and upper limits of the number of hot Jupiters with observable orbital decay.
@@
For the 326 hot Jupiters in Wang et al. (2024) [16], we calculate the theoretical prediction of their decay rate of orbital due to tidal interaction. For the stellar tidal quality factor, we examine three different distributions:
@@
    Results from WASP-12 system [15]: $\log Q^{{}^{\prime}}_{*}=5.14\pm 0.05$.
@@
We adopt the $3\sigma$ errorbars of the observed period change rate as the detection threshold (as shown in Fig. 9 of [16]). Hot Jupiters with calculated $\dot{P}$ values exceeding the detection limit are considered to exhibit detectable orbital decay and we count the expected numbers $N^{\rm detectable}_{\rm decay}$ for the above three $\log Q^{{}^{\prime}}_{*}$ distributions. To consider the effect of uncertainties and obtain the confidence level, we resample the data (i.e., stellar/planetary parameters, $Q^{{}^{\prime}}_{*}$, detection limit) considering the uncertainties for 1,000 times and repeat the above procedure. Then we count the numbers of hot Jupiters with detectable orbital decay $N^{\rm detectable}_{\rm decay}$ by taking different $\log Q^{{}^{\prime}}_{*}$ distributions for each set of resampled data. The 1-$\sigma$ intervals of $N^{\rm detectable}_{\rm decay}$ are taken as their $50\pm 34.1$ percentiles.

### v2

@@ 2.2 The tidal decay of hot Jupiters
Because hot Jupiters are close to their host stars, the tidal interactions can reduce the orbital eccentricity ($e$) of close-in planets [1996Sci...274..954R, 2008ApJ...678.1396J]. Tides can also transfer orbital energy and angular momentum from the hot Jupiters to their host stars, resulting in the spin-up of the stars [2009MNRAS.396.1789P, 2010ApJ...719..602S, 2013ApJ...775L..11M, 2014ApJ...786..139T, 2015A&A...577A..90M]. As a consequence, the semi-major axis ($a$) of hot Jupiters would shrink with age. Part of them may have already passed into the Roche limit and been tidally disrupted by the host star [2008ApJ...678.1396J, 2009ApJ...692L...9L].
In order to quantitatively determine the tidal reduction of the semi-major axis $a$, we adopt the classical equations of equilibrium tidal evolution: [GOLDREICH1966375, 2012MNRAS.423..486L]
@@
where $G$ is the gravitational constant, $R$ and $M$ are the radius and mass, and $Q^{{}^{\prime}}$ is the modified tidal dissipation factor [GOLDREICH1966375]. The subscripts * and $\rm p$ refer to the star and planet, respectively. As can been, due to the tides, $a$ decays with the increase of age. The tidal decay rate is dependent on various parameters (e.g., the initial $a$ and $Q^{{}^{\prime}}_{*}$).
Plenty of previous studies have estimated the stellar tidal quality factor $Q^{{}^{\prime}}_{*}$, using various samples and methodologies. Specifically, Jackson et al. (2008) [2008ApJ...678.1396J], Bonomo et al. (2017) [2017AA...602A.107B], Penev et al. (2018) [2018AJ....155..165P] and Labadie-Bartz et al. (2019) [2019ApJS..240...13L] provided constraints based on the evolution of planetary semi-major axis ($a$), eccentricity ($e$), and stellar spin-up. Whereas, Meibom et al. (2005) [2005ApJ...620..970M] and Milliman et al. (2014) [2014AJ....148...38M] restrict the $Q^{{}^{\prime}}_{*}$ from the tidal interaction between stellar binaries. Collier Cameron & Jardine (2018) [2018MNRAS.476.2542C]estimated the tidal quality factor by reproducing the observed orbital separation distribution, assuming an initially uniform distribution inferred from Doppler surveys. Hamer & schlaufman (2019) [2019AJ....158..190H] obtains a upper limit of $\log Q^{{}^{\prime}}_{*}\lesssim 10^{7}$ by assuming that hot Jupiters were destroyed before the end of the main sequence lifetime of their host stars. Recently, Yee et al. (2020) [2020ApJ... [^8] L...5Y] and Turner et al. (2021) [2021AJ....161...72T] constrained $Q^{{}^{\prime}}{*}$ by analyzing the orbital decay of WASP-12b, a transiting hot Jupiter with a 1.09-day orbit around a late-F star. Additionally, some other studies have detected hints of orbital decay for tens of hot Jupiters (e.g., WASP-4b, 43b) and provided estimations for $Q^{{}^{\prime}}_{*}$ [2020AJ....159..150P, 2022AJ....163..281T]. We have summarized the constraints on the value of $Q^{{}^{\prime}}_{*}$ from the literature in Extended Data Tab. 1.
@@ 2.3 Numerical simulation
    ‘Late’ model. The formation timescale obeys a distribution referring to Fig. 9 of Hamers et al. (2017) [2017MNRAS.464..688H].
@@
For the subsequent tidal evolution, we use a simplified model to perform numerical simulations by setting the orbital eccentricity $e=0$ since the circularization timescale is order of magnitudes less than the in-sprial timescale for hot Jupiters [2008ApJ...678.1396J, 2012MNRAS.423..486L]. That is to say, the conditions for a hot Jupiter’s ‘arrival’ defined in this work refer to it entering within 0.1 AU of its star, with its orbit having been largely circularized. The simulations are carried out by adopting the eighth-order Runge-Kutta method and the precision is set as $10^{-12}$. As shown in Equation S13, the initial conditions consist of the stellar tidal dissipation factor $Q^{{}^{\prime}}_{*}$, the masses and radii of star and planet, and the initial period distribution of hot Jupiters.
@@
    ‘Early’ model: For the planetary mass and radius, we adopt the same distributions as the observed warm/cold Jupiters. For the initial orbital period distribution, we here adopt a distribution inferred from the confirmed planets in Kepler field [2016A&A...587A..64S].
@@
    ‘Late’ model: mass and period as the results of Fig. 8 and 3 of Hamers et al. (2017) [2017MNRAS.464..688H]. The planetary radius are estimated with the planetary mass-radius relation, $\frac{R_{\rm p}}{R_{\rm J}}=1.2\left(\frac{M_{\rm p}}{M_{\rm J}}\right)^{0.27}$ [2016A&A...589A..75M].
@@
In each set of simulation, we put 10,000 hot Jupiters with the aforementioned initial conditions. During the simulations, the hot Jupiters were removed from the system when they fell into the Roche limit, $a_{R}=(R_{p}/0.462)(M_{*}/M_{p})^{1/3}$ [2005Icar..175..248F]. $\log Q^{{}^{\prime}}_{*}$ is set at the interval of 0.1 from 4 to 9. Upon completion of the simulations, the returned information consist of the simulation time, semi-major axis, and the lifetimes of each hot Jupiter.
@@ 3.2 Additional evidence from the obliquity distributions
Stellar obliquity refers to the angle between a planet’s orbital axis and its host star’s spin axis, which can provide important clues to the formation and dynamic evolution of hot Jupiters. Specifically, the initial distribution of stellar obliquity is determined by the formation and evolution of a protoplanetary disk [2010MNRAS.401.1505B, 2011MNRAS.412.2790L]. Subsequently, the stellar obliquity can be further excited by the dynamic mechanisms via interactions caused by planetary and/or binary companions, e.g., planet–planet scattering , stellar/planetary Lidov–Kozai (LK) effects secular chaos. The predicted stellar obliquity distributions of the hot Jupiters via different mechanisms are different (see fig. 22 of [2022PASP..134h2001A]). For example, the majority of hot Jupiter systems caused by the planet-planet scattering mechanism would be aligned but a small fraction of systems would have a diverse range of stellar obliquities [2012ApJ...751..119B]. In contrast, the secular chaos mechanism will result in misaligned systems with obliquities $\lesssim 90^{\circ}$ [2019MNRAS.486.2265T]. Lastly, the tidal interaction could realign the host star’s spin axis with the planet’s orbital axis and thus reduce the stellar obliquity [2022PASP..134h2001A]. Therefore, investigating the distribution of the stellar obliquities of hot Jupiter systems can help to constrain their formation channel and tidal evolution [2022AJ....164...26H, 2023AJ....166..112D].
From observations, typically only the sky-projected stellar obliquity $\lambda$ (the projections onto the plane of the sky) can be mainly measured through the Rossiter–McLaughlin effect [1924ApJ....60...15R, 1924ApJ....60...22M]. To date, the stellar obliquities have been measured for hundreds of hot Jupiters. Previous studies have found observational evidence that many hot Jupiters in misaligned systems are late-arrived within $\lesssim 0.1$ AU long after their parent protoplanetary disks dissipated [2022AJ....164...26H]. Dong & Foreman-Mackey (2023) [2023AJ....166..112D] reported that most hot Jupiter system are aligned and the other misaligned systems have nearly isotropic stellar obliquities with no strong clustering near 90 degrees.
We initialize the sample of hot Jupiters with $\lambda$ measurements based on the Table A of [2022PASP..134h2001A]. We also compare with the TEPCat orbital obliquity catalog (https://www.astro.keele.ac.uk/jkt/tepcat/obli- quity.html) and remove hot Jupiters with the differences in $\lambda$ larger than three times of uncertainties. Then we cross-match with our planetary sample and yield 39 hot Jupiters with $\lambda$ measurements.
As mentioned before, the tidal interaction would reduce the observed $\lambda$. The timescale of tidal-realignment is not well determined yet [2014ARA&A..52..171O]. In the equilibrium tidal theory, the timescale for realignment is comparable to the in-sprial timescale [2022PASP..134h2001A], which is dependent on the stellar mass, radius and planetary mass, radius, orbital semi-major axes. Referring to [2022PASP..134h2001A], we adopt the dimensionless tidal dissipation parameter,
@@
where $M_{*}$ and $R_{*}$ are the stellar mass and radius. $M_{\rm p}$ and $a$ are the planetary mass and semi-major axes. Then we select hot Jupiter systems with long-alignment timescales using the following criterion: $\left(\frac{M_{\rm p}}{M_{*}}\right)^{-2}\left(\frac{a}{R_{*}}\right)^{6}>10^{12}$ or $T_{\rm eff}>6250$ K [2022PASP..134h2001A]. Furthermore, as shown in Fig. 11 of [2022PASP..134h2001A], hot Jupiter systems older than $\sim 4$ Gyr have relatively smaller $\lambda$, thus we also only keep hot Jupiter systems with kinematic ages <4<4 Gyr. We also make Kolmogorov–Smirnov (KS) tests and check the resulted $p$ values to quantify the age evolution.
@@
To further put constraints on the origin of hot Jupiters, we then compare the observed $\lambda$ distribution with the predictions of different formation mechanisms, i.e., planet-planet scattering [2012ApJ...751..119B], resonance crossing [2018MNRAS.480.1402A], secular chaos [2019MNRAS.486.2265T], and planetary Kozai-Lidov effect [2016ApJ...829..132P]. We also conduct an aligned distribution of stellar obliquity $\phi$ as uniformly distributed in the range of $0-20^{\circ}$. Specifically, we first generate 1,000 hot Jupiter systems for each of the above mechanisms with the distributions of stellar obliquity $\phi$. Then we obtain the sky-projected obliquity $\lambda$ distribution from the $\phi$ distribution with the following formula [2023AJ....166..112D]:
@@
where $\theta$ is the azimuthal angle of the stellar spin axis relative to the orbital axis, and $i_{\rm orb}$ is the orbital inclination. Referring to [2023AJ....166..112D], we take $\theta$ and $\cos{i_{\rm orb}}$ as uniformly distributed between 0 and 1. Finally, we perform two-sample KS tests between the $\lambda$ distributions of the observed ‘early-arrived’/‘late-arrived’ hot Jupiter systems and the prediction of the five mechanisms.
The resulted $p$-values are summarized as Extended Data Tab. 3. As can be seen, the $\lambda$ distribution of the ‘early-arrived’ population matches well with the predictions of the planet-planet scattering [2012ApJ...751..119B], the resonance cross [2018MNRAS.480.1402A], and the alignment distribution, but is significantly smaller than the prediction of planetary Kozai-Lidov mechanism [2016ApJ...829..132P] with a KS $p$-value <0.003<0.003. For the ‘late-arrived’ hot Jupiters, their $\lambda$ distribution is consistent with the prediction of the secular chaos mechanism [2019MNRAS.486.2265T] with a KS $p$-value of 0.229, but it deviates from the predictions of all other mechanisms with KS $p$-values<0.05<0.05. The most-favored mechanisms (with largest $p$-values, highlighted in Extended Data Tab. 3) for the ‘early-arrived’ and ‘late-arrived’ hot Jupiters are planet-planet scattering and secular chaos, respectively.
@@ 3.3 Consistency check of stellar metallicity
Stellar metallicity also has important impacts on the formation and evolution of hot Jupiters [2015ApJ...799..229W]. From theory, gentle disk migration is expected to operate in environments with a range of metallicities. Whereas high eccentricity migration via secular chaos occurs in systems with multiple giant planets, which are more likely to form around metal-rich stars born in disks with more solids.
Stellar metallicity is known to be correlated with other stellar properties [2022AJ....163..249C]. Thus we adopt the NearestNeighbors method in scikit-learn to select the nearest neighbor in the stellar mass and $TD/D$ from the ‘early-arrived’ hot Jupiter hosts for every ‘late-arrived’ hot Jupiter host star. As shown in Extended Data Fig. 7, after eliminating the effects of stellar mass, radius and age, the ‘early-arrived’ hot Jupiter host stars have a relative smaller $\rm[Fe/H]$ comparing to the ‘late-arrived’ hot Jupiter hosts. We generate the 10,000 bootstrapped sample from the $\rm[Fe/H]$ of ‘late-arrived’ and ‘early-arrived’ hot Jupiter hosts (after parameter controlling) and calculate their average metallicities $\overline{\rm[Fe/H]}$. The uncertainties are taken as the $50\pm 34.1$ percentile. The $\overline{\rm[Fe/H]}$ are $0.06^{+0.02}_{-0.03}$ and $0.12^{+0.02}_{-0.02}$ dex for the ‘early-arrived’ and ‘late-arrived’ hot Jupiter hosts, respectively. Of the 10,000 set of bootstrapping, the $\overline{\rm[Fe/H]}$ of ‘late-arrived’ hot Jupiter hosts are larger for 9,607 times, supporting our inference that the ‘late-arrived’ hot Jupiters are formed via secular chaos.
@@ 3.4 Consistency check of configurations of hot Jupiter systems
Previous studies have detected hints of the orbital decay of hot Jupiters through monitoring of their transit-time variations (TTVs) [2020ApJ... [^8] L...5Y, 2020AJ....159..150P]. Recently, by combining the high-precision 2 minute cadence transit data provided by TESS with archival data from previous work, Wang et al. (2024) identified 11 candidates with decreasing orbital periods that also passed the leave-one-out cross validation test from 326 hot Jupiters [2024ApJS..270...14W]. Out of the 11 candidates, only the orbital decay of WASP 12 b has been confirmed [2020ApJ... [^8] L...5Y, 2021AJ....161...72T]. For the other 10 candidates, the possibility that mechanisms other than tidal interactions (e.g., apsidal precession [2009ApJ...698.1778R], the Rømer effect [2019AJ....157..217B], and the Applegate effect[2010MNRAS.405.2037W]) may contribute to the variations in orbital periods cannot be ruled out. Therefore, 1 and 11 can be taken as the lower and upper limits of the number of hot Jupiters with observable orbital decay.
@@
For the 326 hot Jupiters in Wang et al. (2024) [2024ApJS..270...14W], we calculate the theoretical prediction of their decay rate of orbital due to tidal interaction. For the stellar tidal quality factor, we examine three different distributions:
@@
    Results from WASP-12 system [2021AJ....161...72T]: $\log Q^{{}^{\prime}}_{*}=5.14\pm 0.05$.
@@
We adopt the $3\sigma$ errorbars of the observed period change rate as the detection threshold (as shown in Fig. 9 of [2024ApJS..270...14W]). Hot Jupiters with calculated $\dot{P}$ values exceeding the detection limit are considered to exhibit detectable orbital decay and we count the expected numbers $N^{\rm detectable}_{\rm decay}$ for the above three $\log Q^{{}^{\prime}}_{*}$ distributions. To consider the effect of uncertainties and obtain the confidence level, we resample the data (i.e., stellar/planetary parameters, $Q^{{}^{\prime}}_{*}$, detection limit) considering the uncertainties for 1,000 times and repeat the above procedure. Then we count the numbers of hot Jupiters with detectable orbital decay $N^{\rm detectable}_{\rm decay}$ by taking different $\log Q^{{}^{\prime}}_{*}$ distributions for each set of resampled data. The 1-$\sigma$ intervals of $N^{\rm detectable}_{\rm decay}$ are taken as their $50\pm 34.1$ percentiles.